# Visual Structures Helps Visual Reasoning: Addressing the Binding Problem in VLMs

### v2

@@ Abstract
Despite progress in Vision-Language Models (VLMs), their capacity for visual reasoning is often limited by the binding problem: the failure to reliably associate perceptual features with their correct visual referents. This limitation underlies persistent errors in tasks such as counting, visual search, scene description, and spatial relationship understanding. A key factor is that current VLMs process visual features largely in parallel, lacking mechanisms for spatially grounded, serial attention. This paper introduces a simple yet effective intervention: augmenting visual inputs with low-level spatial structures (e.g., horizontal lines) and pairing this with a textual prompt that encourages sequential, spatially-aware parsing. We empirically demonstrate substantial performance improvements across core visual reasoning tasks. Specifically, our method improves GPT-4o visual search accuracy by 25.00%, increases counting accuracy by 26.83%, reduces edit distance error in scene description by 0.32, and enhances performance on spatial relationship tasks by 9.50% on a a 2D synthetic dataset. Furthermore, we find that the visual modification is essential for these gains; purely textual strategies, including Chain-of-Thought prompting, are insufficient and can even degrade performance. Our method enhances binding only with a single-query inference, underscoring the importance of visual input design over purely linguistically-based approaches. These findings suggest that low-level visual structuring is a powerful and underexplored direction for improving compositional visual reasoning and could serve as a general strategy for enhancing VLM performance on spatially grounded tasks.
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable progress, matching or even surpassing human performance across a range of complex tasks [1, 2, 3, 4, 5]. Despite significant advancements, Vision-Language Models (VLMs) still lag behind LLMs in core reasoning capabilities. While LLMs excel at symbolic and abstract reasoning, VLMs struggle with essential aspects of visual understanding. They frequently miscount objects in cluttered or overlapping scenes [6, 7], perform poorly in spatial reasoning (e.g., identifying “left of” or “above”) [8], and often fail to accurately locate or recognize salient targets during visual search [9]. Moreover, compositional and relational reasoning, where understanding critically depends on structured relationships between objects, remains a major challenge [6]. These limitations indicate that VLMs, despite their power, lack key mechanisms for structured visual reasoning. Consequently, their performance in these areas does not yet match that of language-only models in comparable text-based reasoning tasks.
@@
The limitations in visual reasoning tasks can largely be attributed to a fundamental challenge known as the binding problem [9]. Originating from cognitive science and neuroscience, the binding problem refers to the difficulty of correctly associating visual features (like shape and color) and spatial properties (such as location and orientation) with the right objects in a scene. It arises when a system must handle multiple entities simultaneously, leading to interference and misattribution of features [10, 11, 12]. In the context of VLMs, the binding problem manifests as errors in compositional understanding—models may recognize individual objects correctly but confuse or conflate their attributes. These issues are particularly prevalent in complex, cluttered scenes and persist across a range of model architectures and benchmarks. As a result, reliable multi-object reasoning and generalization of visual concepts to novel settings remain difficult for current VLMs.
@@
Figure 1: Overview of the proposed method. The input image is augmented with simple spatial scaffolding using 3 horizontal lines to impose a low-level visual structure. A corresponding textual prompt (“Scan the image sequentially based on horizontal lines”) is appended to encourage the model to adopt a spatially guided, sequential parsing strategy.
@@
Inspired by the neuroscientific studies on human scene perception, in this work, we propose a novel approach to improve VLM binding performance. Our approach augments visual inputs with low-level spatial structures, such as horizontal or grid lines, complemented by appending a textual instruction to the input prompt (See Figure 1 as an example). The motivation behind our strategy is to encourage sequential processing in VLMs guided by these interventions in the inputs. Our method can be viewed as a visual analogue of Chain-of-Thought (CoT) prompting[13] for LLMs, as it injects an expert-designed inductive bias directly into the visual input to implicitly decompose the problem and guide the model’s reasoning.
@@
Despite its simplicity, our method yields significant improvements in performance. Moreover, we empirically show that both visual and textual components of our method are necessary and complementary. In particular, textual prompts for sequential scanning alone prove insufficient to resolve binding failures, and even attempts to elicit more structured reasoning through complex prompting strategies, such as Chain-of-Thought (CoT), can paradoxically degrade performance on these visual tasks [14]. In contrast, the inclusion of visual cues stimulates more structured internal processing in VLMs and promotes spatial parsing mechanisms, enhancing the model’s ability to bind features accurately. These findings highlight the potential of targeted visual input modifications to improve graphical reasoning and compositional understanding in VLMs. In summary, the main contributions of this paper are:
@@
  • We propose a novel method that combines explicit visual scaffolding (e.g., horizontal lines) with targeted textual prompts to improve feature-object binding in VLMs.
@@
  • Through extensive experiments, we demonstrate that our method significantly enhances VLM performance across a range of core visual reasoning tasks, including visual search, counting, scene description, and spatial relationship understanding. For example, on a 2D scene understanding task using GPT-4o, we improve the accuracy of visual search by 25.00%, counting by 26.83%, and spatial relationships by 9.50%, while also reducing the edit distance error for the scene description task by 0.32.
@@
  • The proposed intervention achieves these improvements efficiently, operating within a single query and incurring negligible computational overhead. This efficiency distinguishes our method from multi-query or existing agentic approaches.
@@ 2 Related Work
LLM and VLM Reasoning. While early LLMs were regarded as next-token predictors with little reasoning ability [15, 16], recent work has begun to extensively challenge this problem [17, 18, 19, 20]. Today’s state-of-the-art models deliver high performance across many tasks, including graduate-level question answering [21] and competitive programming. The improved reasoning abilities of LLMs have naturally sparked interest in reasoning for VLMs. Initial approaches include Visual Chain-of-Thought [22], knowledge graph integration [23], and tree search [24]. Although these techniques show promise, VLMs still struggle with tasks such as counting, visual search, scene description, and spatial reasoning, as demonstrated by benchmarks like EMMA [14] and SPACE [25]. Our work aims to incorporate an explicit reasoning pathway into the visual component of the model to improve its performance on these challenging tasks.
Binding Problem. The binding problem offers a key explanation for the challenges faced by state-of-the-art VLMs in associating different features of objects, such as shape and location.  [26, 27]. Some studies suggest that performance issues in tasks like counting, visual search, and scene description can be attributed to binding problems in models [9]. Recent neuroscience research has shown that grid-based frameworks enhance visual recognition memory [28] and improve face recognition performance [29]. Additionally, humans reduce interference by detecting individual objects iteratively [30, 10], and grid structures facilitate movement-based object recognition [31], thus providing insight into how the human brain can deal with the binding problem. Inspired by these insights, we use horizontal lines in images to enhance model reasoning, both with synthetic images and real-world datasets, such as the Everything Counts benchmark [32].
Agentic VLMs. An emerging line of research treats VLMs as autonomous tool users that can invoke external modules or produce executable artefacts to compensate for perceptual or reasoning gaps. LVLM-COUNT [33] decomposes enumerative queries into sub-counts solved by a specialized counting utility, yielding gains on counting benchmarks. Visual Sketchpad [34] endows multimodal LMs with a drawable canvas, letting them sketch auxiliary lines and marks. Code execution agents such as ViperGPT [35] translate natural-language questions into Python scripts that orchestrates off-the-shelf vision tools, achieving state-of-the-art results on compositional visual queries. While Toolformer [36] shows that language models can self-supervise the decision of when and how to call external APIs. Although these approaches boost task performance, they do not enrich the models’ intrinsic reasoning capabilities. In contrast, our work embeds an explicit reasoning pathway directly into the visual input to overcome these limitations.
@@ 3 Proposed Method
The binding problem—a fundamental challenge in cognitive science and neuroscience—concerns how perceptual systems correctly associate features like color, shape, and location with the right objects in complex visual scenes [10, 12]. When multiple objects share representational resources, the system can produce illusory conjunctions, i.e., erroneous confusion of features from different objects (e.g., mistakenly perceiving a red square when shown a red circle and a green square) [11]. These misbindings reveal the inherent difficulty of forming coherent, object-specific representations from a distributed set of visual information and analyzing them offers a valuable framework for assessing the capability and reliability of visual reasoning systems across diverse tasks.
Neuroscientific studies show that human visual system operates in two distinct modes:  [^1] Also called System 1 and System 2 in the context of visual reasoning [37]. Parallel processing is fast but imprecise, particularly in cluttered scenes where binding demands are high. When given sufficient time budget, humans perform a more accurate feature binding by transitioning from the parallel analysis to the sequential attention. In contrast, current VLMs process inputs predominantly in parallel. This parallel feature extraction is a double-edged sword; while it enables compositional representation learning—supporting flexibility and generalization—it inherently creates the risk for interference and binding errors, particularly if feature-to-object linkages are not meticulously managed [27].
To address the binding problem in VLMs, we propose a lightweight, model-agnostic approach that enhances structured visual reasoning through minimal visual scaffolding and spatially grounded prompting. Inspired by cognitive science findings on serial attention and spatial structuring, our method guides VLMs to process visual information region-by-region and reduces interference effects from parallel processing, a known contributor to binding errors in human and machine vision. This approach introduces negligible computational overhead, requires no task-specific fine-tuning or multi-query visual processing, and comprises two main components, which are illustrated with an example in Figure 1 and explained below:
@@
We augment each input image with $n$ equidistant horizontal lines, partitioning it into $n+1$ segments numbered sequentially from top to bottom ($1$ through $n+1$). These lines serve as visual anchors, promoting localized attention within each region to reduce cross-object interference and improve feature-object binding. Unlike dense grid-based methods that may obscure content, our minimalist design preserves clarity while providing sufficient spatial guidance. The value of $n$ is set to $3$ in our experiments.
@@
To align the model’s attention with the visual scaffold, we prepend a fixed instruction, specifically: “Scan the image sequentially based on horizontal lines exists in the image.” This prompt guides the model to adopt a structured, row-wise processing strategy, encouraging a systematic evaluation of the image content. For task-specific adaptation (e.g., counting or spatial reasoning), we augment the base prompt with additional instructions (see Appendix D).
The use of manually added artifacts and the emphasis on sequential, region-based processing may also align with the discussion of feature entropy in [9]. By partitioning the image into distinct segments, our method encourages focused processing on smaller subsets of objects. This can simplify the distinction between objects within each segment by reducing the concurrent feature load, an effect pertinent to managing feature diversity (or entropy) as detailed in [9]. Such simplification is a key factor contributing to improved visual reasoning, particularly in counting.
@@ 4 Experiments
Figure 2: A brief summery of tasks with one example of synthetic data along the specific prompt for each task. (a) Visual Search, (b) Spatial Relationship, (c) Counting, and (d) Spatial Relationship.
@@ 4.1 Setup
To evaluate our method, we use both synthetic and natural datasets. A Binding Problem Generator [9] produces synthetic data that can be controlled in two dimensions and three dimensions and can incorporate a variable number of objects. Additionally, we benchmark on two real-world tasks: Learning To Count Everythig [32] and the Spatial Reasoning [38] datasets.
@@
We evaluated our proposed method across a carefully chosen mix of closed-source and open-source VLMs. Specifically, we include two leading closed-source systems—OpenAI’s GPT-4o [39] and Anthropic’s Claude3.5-sonnet [40]—selected for their cutting-edge capabilities and solid performance on multimodal benchmarks. For GPT-4o, we evaluate both our method and Chain-of-Thought (CoT) prompting [13] to enable a direct comparison within the same architecture. Additionally, we benchmark two state-of-the-art open-source models: Qwen2.5-VL-7B-Instruct [41] and LLaMa4-scout-17b-16e-Instruct [42]. Beyond these baselines, we assess the benefits of targeted fine-tuning by incorporating Mulbary [24], a Qwen2.5VL-7B variant refined for enhanced multimodal reasoning. This breadth of model selection allows us to rigorously assess the generality and robustness of our approach across diverse architectures, training scales, and access constraints.
@@
In addition to the main experiments reported in the following subsections, a set of ablation studies is provided in Appendix A.
4.2 Visual Search
Table 1: Harmonic mean comparison for the Visual Search task, evaluating our method against base models across GPT-4o, Claude3.5-sonnet, LLaMa4, and Qwen2.5-VL on 2D and 3D datasets with varying numbers of objects.

### v4

# Visual Structures Help Visual Reasoning: Addressing the Binding Problem in LVLMs

@@ Abstract
Despite progress in Large Vision-Language Models (LVLMs), their capacity for visual reasoning is often limited by the binding problem: the failure to reliably associate perceptual features with their correct visual referents. This limitation underlies persistent errors in tasks such as counting, visual search, scene description, and spatial relationship understanding. A key factor is that current LVLMs process visual features largely in parallel, lacking mechanisms for spatially grounded, serial attention. This paper introduces Visual Input Structure for Enhanced Reasoning (VISER), a simple, effective method that augments visual inputs with low-level spatial structures and pairs them with a textual prompt that encourages sequential, spatially-aware parsing. We empirically demonstrate substantial performance improvements across core visual reasoning tasks, using only a single-query inference. Specifically, VISER improves GPT-4o performance on visual search, counting, and spatial relationship tasks by 25.0%, 26.8%, and 9.5%, respectively, and reduces edit distance error in scene description by 0.32 on 2D datasets. Furthermore, we find that the visual modification is essential for these gains; purely textual strategies, including Chain-of-Thought prompting, are insufficient and can even degrade performance. VISER underscores the importance of visual input design over purely linguistically based reasoning strategies and suggests that visual structuring is a powerful and general approach for enhancing compositional and spatial reasoning in LVLMs. Webpage is available at https://sharif-ml-lab.github.io/VISER/.
@@ 1 Introduction
Large Language Models (LLMs) have recently achieved remarkable progress, matching or even surpassing human performance across a range of complex tasks [1, 2, 3, 4, 5]. Despite significant advancements, Large Vision-Language Models (LVLMs) still lag behind LLMs in core reasoning capabilities. While LLMs show strong symbolic and abstract reasoning, LVLMs struggle with essential aspects of visual understanding. They frequently miscount objects in cluttered or overlapping scenes [6, 7], perform poorly in spatial reasoning (e.g., identifying “left of” or “above”) [8], and often fail to accurately locate or recognize salient targets during visual search [9]. Moreover, compositional and relational reasoning, where understanding critically depends on structured relationships between objects, remains a major challenge [6]. These limitations indicate that LVLMs, despite their power, lack key mechanisms for structured visual reasoning. Consequently, their performance in these areas does not yet match that of language-only models in comparable text-based reasoning tasks.
@@
The limitations in visual reasoning tasks can largely be attributed to a fundamental challenge known as the binding problem [9]. Originating from cognitive science and neuroscience, the binding problem refers to the difficulty of correctly associating visual features (like shape and color) and spatial properties (such as location and orientation) with the right objects in a scene. It arises when a system must handle multiple entities simultaneously, leading to interference and misattribution of features [10, 11, 12]. In the context of LVLMs, the binding problem manifests as errors in compositional understanding—models may recognize individual objects correctly but confuse or conflate their attributes. These issues are particularly prevalent in complex, cluttered scenes and persist across a range of model architectures and benchmarks. As a result, reliable multi-object reasoning and generalization of visual concepts to novel settings remain difficult for current LVLMs.
@@
Figure 1: The input image is augmented with low-level visual structure using three horizontal lines, optionally accompanied by row annotations on the left side of the image. A corresponding textual prompt (“Scan the image sequentially based on horizontal lines”) is appended to encourage the model to adopt a spatially guided, sequential parsing strategy.
@@
Inspired by neuroscientific studies on human scene perception, we propose Visual Input Structure for Enhanced Reasoning (VISER), a novel approach to improve LVLM binding performance. Our approach augments visual inputs with low-level spatial structures, such as horizontal or grid lines, complemented by appending a textual instruction to the input prompt (see Figure 1 as an example). The motivation behind our strategy is to encourage sequential processing in LVLMs guided by these structures in the inputs. Our method can be viewed as a visual analogue of Chain-of-Thought (CoT) prompting[13] for LLMs, as it injects an inductive bias directly into the visual input to implicitly decompose the problem and guide the model’s reasoning.
@@
Despite its simplicity, VISER yields significant performance improvements. The method is general and input-agnostic, requiring no additional processing for specialized image manipulation. We empirically show that both the visual structure and the textual part of our method are necessary and complementary. In particular, textual prompts for sequential scanning alone prove insufficient to resolve binding failures, and even attempts to elicit more structured reasoning through complex prompting strategies, such as CoT, can paradoxically degrade performance on these visual tasks [14]. In contrast, the inclusion of visual cues stimulates more structured internal processing in LVLMs and promotes spatial parsing mechanisms, enhancing the model’s ability to bind features accurately. These findings highlight the potential of targeted visual input modifications to improve graphical reasoning and compositional understanding in LVLMs. In summary, the main contributions of this paper are:
@@
  • We propose VISER, a novel method that combines explicit visual scaffolding (e.g., horizontal lines) with targeted textual prompts to improve feature-object binding in LVLMs.
@@
  • Through extensive experiments, we demonstrate that our method significantly enhances LVLM performance across a range of core visual reasoning tasks, including visual search, counting, scene description, and spatial relationship understanding. For example, on 2D benchmark, VISER improves GPT-4o performance on visual search, counting, and spatial reasoning by 25.0%, 26.8%, and 9.5% (absolute), respectively, and reduces the edit-distance error in scene description by 0.32.
@@
  • The proposed approach achieves these improvements efficiently, operating within a single query and incurring negligible computational overhead. Its effectiveness, simplicity, and generality (both task- and model-agnostic) distinguish our method from multi-query or agentic approaches that rely on additional processing or tool use.
@@ 2 Related Work
LLM and LVLM Reasoning. While early LLMs were regarded as next-token predictors with little reasoning ability [15, 16], recent work has begun to extensively challenge this problem [17, 18, 19, 20]. Today’s state-of-the-art models deliver high performance across many tasks, including graduate-level question answering [21] and competitive programming. The improved reasoning abilities of LLMs have naturally sparked interest in reasoning for LVLMs. Initial approaches include Visual Chain-of-Thought [22], knowledge graph integration [23], and tree search [24]. Although these techniques show promise, LVLMs still struggle with tasks such as counting, visual search, scene description, and spatial reasoning, as demonstrated by benchmarks like EMMA [14] and SPACE [25]. Our work aims to incorporate an explicit reasoning pathway into the visual component of the model to improve its performance on these challenging tasks.
Binding Problem. The binding problem provides a key explanation for the limitations of state-of-the-art LVLMs in reliably linking visual features to their corresponding objects [26, 27]. Prior work attributes performance issues in tasks such as counting, visual search, and scene description to failures in this feature–object binding in LVLMs [9]. Recent neuroscience research has shown that grid-based frameworks enhance visual recognition memory [28] and improve face recognition performance [29]. Additionally, humans reduce interference by detecting individual objects iteratively [30, 10], and grid structures facilitate movement-based object recognition [31], thus providing insight into how the human brain can deal with the binding problem. Inspired by these insights, we use horizontal lines in images to enhance model reasoning, both with synthetic images and real-world datasets, such as the Learning To Count Everything benchmark [32].
Agentic LVLMs. An emerging line of research treats LVLMs as autonomous tool users that can invoke external modules or produce executable artefacts to compensate for perceptual or reasoning gaps. LVLM-COUNT [33] decomposes enumerative queries into sub-counts solved by a specialized counting utility, yielding gains on counting benchmarks. Visual Sketchpad [34] endows multimodal LMs with a drawable canvas, letting them sketch auxiliary lines and marks. Code execution agents such as ViperGPT [35] translate natural-language questions into Python scripts that orchestrates off-the-shelf vision tools, achieving state-of-the-art results on compositional visual queries. Toolformer [36] shows that language models can self-supervise decisions about when and how to call external APIs. Although such tool-using approaches improve task performance, they do not necessarily enrich models’ intrinsic reasoning capabilities. In contrast, our work embeds an explicit reasoning pathway directly into the visual input to overcome these limitations.
@@ 3 Proposed Method
The binding problem, a fundamental challenge in cognitive science and neuroscience, concerns how perceptual systems correctly associate features like color, shape, and location with the right objects in complex visual scenes [10, 12]. When multiple objects share representational resources, the system can produce illusory conjunctions, i.e., erroneous confusion of features from different objects (e.g., mistakenly perceiving a red square when shown a red circle and a green square) [11]. These misbindings reveal the difficulty of forming coherent, object-specific representations from distributed visual inputs; their analysis provides a useful framework for evaluating visual reasoning systems.
Neuroscientific studies indicate that the human visual system operates in two distinct modes: a fast, imprecise parallel processing (often likened to System 1 processing) and a more accurate sequential attention (System 2 reasoning) [37]. While parallel processing enables rapid scene understanding, it is prone to errors, particularly in cluttered environments where the demands for feature binding are high. Given sufficient time, humans transition to sequential attention, which allows for more meticulous feature binding and reduces errors. In contrast, current LVLMs predominantly process visual inputs largely in parallel. This approach, while beneficial for compositional representation learning and generalization, inherently risks interference and binding errors if feature-to-object linkages are not precisely managed [27].
To address the binding problem in LVLMs, we propose a lightweight, model-agnostic approach that enhances structured visual reasoning through minimal visual scaffolding and spatially grounded prompting. Inspired by cognitive science findings on serial attention and spatial structuring, our method guides LVLMs to process visual information region-by-region and reduces interference effects from parallel processing, a known contributor to binding errors in human and machine vision. This approach introduces negligible computational overhead, requires no task-specific fine-tuning or multi-query visual processing, and comprises two main components, which are illustrated with an example in Figure 1 and explained below:
@@
We augment each input image with $n$ equidistant horizontal lines, creating $n+1$ horizontal segments. Each segment is annotated with a numerical label ($1$ through $n+1$) placed sequentially from top to bottom. These lines serve as visual anchors, promoting localized attention within each region to reduce cross-object interference and improve feature-object binding. Unlike dense grid-based methods that may obscure content, our minimalist design preserves clarity while providing sufficient spatial guidance. The value of $n$ is set to $3$ in our experiments.
@@
To align the model’s attention with the visual scaffold, we prepend a fixed instruction, specifically: “Scan the image sequentially based on horizontal lines exists in the image.” This prompt guides the model to adopt a structured, row-wise processing strategy, encouraging a systematic evaluation of the image content. For task-specific adaptation (e.g., counting or spatial reasoning), we augment the base prompt with additional instructions (see Appendix C).
The use of manually added artifacts and the emphasis on sequential, region-based processing may also align with the discussion of feature entropy in [9]. By partitioning the image into distinct segments, our method encourages focused processing on smaller subsets of objects. This can simplify the distinction between objects within each segment by reducing the concurrent feature load, an effect pertinent to managing feature diversity (or entropy) as detailed therein. Such simplification is a key factor contributing to improved visual reasoning, particularly in counting.
@@ 4 Experiments
Figure 2: A brief summary of tasks with one example of synthetic data along the specific prompt for each task. (a) Visual Search, (b) Scene Description, (c) Counting, and (d) Spatial Relationship.
@@ 4.1 Setup
To evaluate VISER, we use both synthetic and natural datasets. A Binding Problem Generator [9] produces synthetic data that can be controlled in two and three dimensions and can incorporate a variable number of objects. Additionally, we benchmark on two real-world tasks: Learning To Count Everything [32] and the Spatial Reasoning [38] datasets.
@@
We evaluated our proposed method across a carefully chosen mix of closed-source and open-source LVLMs. Specifically, we include two leading closed-source systems—OpenAI’s GPT-4o [39] and Anthropic’s Claude3.5-sonnet [40]—selected for their cutting-edge capabilities and solid performance on multimodal benchmarks. For GPT-4o, we evaluate both our method and Chain-of-Thought (CoT) prompting [13] to enable a direct comparison within the same architecture. Additionally, we benchmark two state-of-the-art open-source models: Qwen2.5-VL-7B-Instruct [41] and LLaMa4-scout-17b-16e-Instruct [42]. Beyond these baselines, we assess the benefits of targeted fine-tuning by incorporating Mulberry [24], a Qwen2.5VL-7B variant refined for enhanced multimodal reasoning. This breadth of model selection allows us to rigorously assess the generality and robustness of our approach across diverse architectures, training scales, and access constraints.
@@
In addition to the main experiments, ablation studies are provided in Appendix A, where we isolate the effect of each component of VISER and evaluate various hyperparameters across different models. Specifically, we assess variants that use only the visual scaffolding or only the sequential prompt (see Appendix A.1), as well as different numbers of structural lines. Moreover, we investigate the impact of other design choices, such as the number and thickness of lines (Appendix A.2). These analyses confirm that both components contribute to the overall performance gains and that VISER remains effective across a reasonable range of hyperparameter settings. Finally, in Appendix B.3, we show the significance of the proposed method by applying a binomial sign test on the results.
4.2 Counting
Counting involves determining the number of specific objects within a scene. Although it may appear simple, accurate enumeration—especially when multiple object types or attribute variations are involved—relies heavily on effective feature binding. Each object must be individuated, its defining features correctly associated, and then counted as a distinct instance. Failures in binding can lead to omissions, double-counting, or misclassification. Similar to human rapid estimation under time constraints, LVLMs show capacity limits in counting. Performance often worsens as object numbers grow, or when high object similarity (low feature heterogeneity) increases the interference and binding errors.
Table 1: Performance comparison on the counting task using models GPT-4o, Claude3.5-sonnet, LLaMa4, and Qwen2.5-VL, evaluated with accuracy (%) . The comparison is conducted across 2D and 3D datasets with varying numbers of objects in the scene, as well as a natural image dataset. Results are shown for base models and VISER.